please answer EVERY question for an upvote ?? + good review!!

Find the antiderivative. 1. / (2 + z ² )d 2. / sin(x) da 3. / sin(4x)dex 4. /le+e->)de 5. / (r sin(x)de 6. / 12V 22 – 5}dt
7. 8. sin (cos ®) de 9. :da 10 dr In z
Answer #1

Ag! 0 [(a+2) da = tegas + +C {1x {1x?d7=xt nt) 3 @ Sin(x) dx --CA(x) a + c (sin(4x) dx = -cos (42) S sinar.da zlogaz + C CL 4J 7²-5 0 0 싸 우 ~5 dt da 보일 이와는 20d 2 { ay = x++) C 노 0) 2 치 ſt at It. 中C 뼈는 ) g () (14.c, fit it > i7-5 3 +C Put it 2 x 2 -5chince 201-235423 (1-7) 11-x) + c 8 f sind (6000).do 3 NIFT Let Cos oct at - Sino -dt = Sino.do z do 3.dt U ot TH put t=cosoWAL 10 Szent dal alna Let Inx=f dt dx kp. Tato 2 x a Z side ent t Put tzena / henne = xp si An (10 (x)) to хч Х.